# RingMoE: Mixture-of-Modality-Experts Multi-Modal Foundation Models for Universal Remote Sensing Image Interpretation

### v1

@@ 1 Introduction
Foundation models have recently emerged as transformative in AI, leveraging self-supervised learning [5] to extract generalized, task-agnostic representations from large-scale unlabeled data. Compared to smaller, task-specific models tailored for individual scenarios, these “one-for-all” general foundation models demonstrate superior performance and generalization across diverse interpretation tasks, particularly in data-scarce environments. Notable examples such as ChatGPT [6] in NLP, SiMIM [7] and MAE [8] in CV, and emerging models in specialized fields like medical imaging [9, 10] and meteorology [11, 12] have redefined benchmarks in their respective domains.
@@
(1) Unified Multi-modal Perception. Building a complete Earth observation system usually requires multiple sensors (e.g., Opt, MS, and SAR) working in tandem to capture Earth’s surface properties. However, most existing RSFMs only serve single or narrow modalities (see Tab.I), e.g. RingMo [16] focus on Opt and DINO-MM [39] on Opt and SAR. Indeed, such modality-specific tailored models hinder the enormous potential of exploring RS multi-modal data. Numerous works have revealed that although heterogeneous sensors have different imaging modes (active imaging for SAR and passive imaging for Opt/MS) and non-uniform technological regimes (high-resolution imaging, wide-area observation, and dynamic monitoring, etc), the information from different modalities can complement each other  [40], e.g., the fusion of Opt and infrared (MS) can improve the imaging quality during nighttime and bad weather conditions [41]. At this point, combining these modalities to construct a unified RS cross-modal representation can fully unleash the potential of multi-modal data and facilitate a better understanding of RS images.
Meanwhile, most existing SAR-based RSFMs treat SAR amplitude data (namely SAR-L2 in this paper) as ordinary CV images for pre-training [23, 24], overlooking the rich spectral information (amplitude and phase) and the unique physical imaging characteristics of SAR sensors. However, analyzing both the amplitude and phase components of radar signals provides crucial insights into surface structures, such as object contours and elevation [42, 43]. For instance, the phase of the non-diagonal elements in the covariance/coherence matrix aids in distinguishing different types of scatterers. Therefore, leveraging the full potential of complex-valued SAR data (namely SAR-L1) is essential for more comprehensive surface object analysis.
@@ 2.1 Remote Sensing Foundation Models (RSFMs)
Self-supervised learning [5, 8, 7] in the CV field has significantly contributed to the flourishing of remote sensing foundation models (RSFMs) [27, 26, 37, 18, 17, 20, 38, 24, 16, 23, 39], which can leverage the vast amount of unlabeled RS images to efficiently capture the generic feature representation. Mainstream RSFMs mainly employ masked image modeling (MIM) and contrastive learning (CL) for pre-training. (1) MIM-based RSFM, which learns image representations by reconstructing masked patches. RingMo [16] optimizes the masking strategy in MIM to focus on dense and small objects. SpectralGPT [19] designs a 3D tensor-shaped spatial-spectral mask strategy to comprehensively capture locally spatial-spectral features and spectral sequential information. Scale-MAE [17] reconstructs low-frequency and high-frequency information of the masked image at different scales to improve multi-scale perception. (2) CL-based RSFM, which optimizes the model by closing the distance between positive samples and widening the distance between negative samples. GASSL [50] takes time-series views as positive sample pairs to neglect temporal variation and focus on more spatial variation. SeCo [51] constructs seasonal-based multi-augmentation contrast loss to perceive short- and long-term temporal changes.
@@ 2.3 Mixture-of-Experts (MoE)
Experience with deep learning suggests that increasing model capacity and dataset scale tends to improve performance [44]. However, it must be acknowledged that scaling models to mega-scale also requires extremely high computational costs. Mixture-of-Experts (MoE) architecture [53, 45, 47, 48] is a promising solution to keep the computational cost acceptable while scaling up the model. With the sparsity of the MoE, the NLP community has expanded language models to a considerable scale [48, 45, 54, 55]. Switch Transformer [48] optimizes the MoE routing pattern and builds a model with up to a trillion parameters. ECR [54] proposes a heterogeneous Mixture-of-Experts employing an expert choice method to further improve the training efficiency. MoA [55] proposes a new architecture that combines multiple attention and the MoE technique. DeepseekMoE [45] constructs fine-grained expert segmentation and shared expert isolation to increase expert specialization and model parameter efficiency (see Appendix A for details on typical MoE architecture). In addition, the researchers have also applied the MoE technique to the CV field [34, 54, 56, 47], e.g, V-MoE first applies the MoE to the CV field and proposes a 15B parameter model, which shows the potential of scale vision models. $\mathbf{M}^{3}$iT [56] and Mod-Squad [47] also design efficient multi-task learning models with the MoE technology. In this paper, we apply the MoE technique to remote sensing and construct a 14.7B parameter RSFM. Notably, for maximizing the parameter potential, we propose modal-specialized, collaborative, and shared experts to rationally divide up the learning of RS multi-modal knowledge.
Figure 5: Overview of the proposed RingMoE framework. Given the multi-modal inputs, the random masking operation is performed to generate visible and masked patches, i.e. tokens. The RingMoE encoder incorporates the RMoE layer (i.e., a novel hierarchical Mixture-of-Experts structure) to replace the standard FFN, effectively capturing both inter-modal commonalities and intra-modal specializations in RS. Finally, the latent representations from different modalities are decoded by modal-specific decoders to reconstruct the original targets.
@@ 3 Proposed RingMoE
This section outlines the pre-training, architecture, efficient deployment, and pre-training dataset. Sec.3.1 provides an overview of the RingMoE pre-training process, followed by Sec.3.2 and Sec.3.3, which detail the feature encoder and modal-specific decoder. Sec.3.4 discusses the expert pruning strategy for effective deployment on various downstream tasks. Finally, Sec.3.5 introduces the multi-modal dataset for pre-training.
@@ 3.2 Joint Encoder across Modalities and Experts for Multi-Modal Representation
RingMoE utilizes Swin Transformer v2 [33] as its image encoder, replacing the standard Feed-Forward Network (FFN) layer with the MoE layer to enhance model capacity and improve adaptability to multi-modal RS data (detailed in Appendix B.2). As illustrated in Fig.5, the image tokens from different modalities $x_{m}$, are first transformed into feature embeddings by modal-specific linear projections $E_{e}$. Then, after embedding the position encoding $E_{pos}$, all the feature embeddings $z_{m}$ from different modalities are fed into the RingMoE $f_{\theta}$ to learn the latent representations $z_{m}$, i.e., $z_{m}=f_{\theta}\left(E_{e}x_{m}+E_{pos}\right)$. Notably, the encoder $f_{\theta}$ consists of $L$ blocks, each of which is built from the Multi-head attention (MHA) and RMoE Layer.
@@ 3.2.1 RMoE Layer
To further enhance model capacity and fully exploit the potential of RSFMs, we propose replacing each FFN layer in the Swin Transformer with the MoE structure (see Appendix A for typical MoE implementations). However, singularly sharing a group of experts across different RS modalities may lead to interference between unrelated or even conflicting modalities, causing confusion or degradation in expert knowledge. In other words, while sharing experts across certain modalities can be beneficial, it can also be detrimental when applied to modalities with distinct imaging characteristics, such as Opt and complex-valued SAR-L1 data. This challenge highlights the inherent difficulty of training a unified multi-modal RSFM.
@@
Shared expert. The shared expert consolidates global knowledge across all modalities, bypassing the complexity of a routing network by directly processing feature tokens from all inputs. This design reduces reliance on additional parameters in the modal-specialized and collaborative experts, enhancing efficiency while maintaining strong performance. By integrating cross-modal information, the shared expert reduces redundancy and ensures robust generalization.
@@ 3.2.2 Optimization of the RMoE layer
$f_{k}=\frac{1}{N}\sum_{i=1}^{N}\mathds{1}{(}{\mathrm{Token}\ x_{m;i}\ \mathrm{selects}\ \mathrm{Expert}\ k})$ (6)
@@
Where $N_{C}$ denotes the number of cooperative experts. $f_{k}$ denotes the probability of tokens dispatched to expert $k$ (see Eq.6), where $N$ denotes the number of the input tokens, $\mathds{1}{(}\cdot)$ is used to determine whether the $i^{th}$ input token $x_{m}$ is assigned to expert $k$ after passing through the routing network $G^{C}$ of the cooperating experts. $H\left(\cdot\right)$ denote the routing network $G^{C}\left(\cdot\right)$ without top-K function.
@@ 3.3 Modal-Specific Decoders for Physics-Informed Self-Supervised Target Reconstruction
For Opt, MS, and SAR-L2 images, we reconstruct the original pixel values to preserve spatial and spectral integrity. For fully polarimetric SAR-L1 data, which consists of four polarization modes with complex-valued representations (each mode containing real and imaginary components), we adopt a physics-informed reconstruction strategy. Instead of reconstructing raw pixel values, we leverage the principle of power conservation before and after polarization decomposition [49], reconstructing the image power as: $t_{SAR_{L1}}=\mathrm{Power}\left(I_{SAR_{L1}}\right)=\left|I_{SAR_{L1}}^{HH}\right|^{2}+\left|I_{SAR_{L1}}^{HV}\right|^{2}+\left|I_{SAR_{L1}}^{VH}\right|^{2}+\left|I_{SAR_{L1}}^{VV}\right|^{2}$. Here, HH, HV, VH, and VV represent the four polarization modes. This approach ensures that the reconstruction task remains physically meaningful, enhancing the model’s ability to capture SAR polarimetric characteristics.
@@
Where $\hat{x}_{m}$ and $t_{m}$ denote the reconstructed target and the original target from modal $m$, respectively. $M\left(\cdot\right)$ denotes the set of masked pixels and $\Omega\left(\cdot\right)$ denotes the number of elements.
@@
Where $\alpha$ is employed to balance the contribution of two losses.
Figure 7: Details of Expert pruning. After pre-training, RingMoE supports modular decomposition into single-modal or multi-modal sub-models of varying scales via pruning, enabling flexible adaptation to downstream tasks. RingMoE can be compressed into modality-specific (e.g., 6.5B) or lightweight (e.g., 1B) models through expert pruning, retaining adaptability for resource-constrained environments with minimal performance degradation.
@@ 3.4 Dynamic Expert Pruning for Efficient Deployment across Various Computational Resources
Following large-scale pre-training, RingMoE ($f_{\theta}$) serves as a universal image encoder for downstream tasks, while the modal-specific decoders ($g_{\phi}$) are discarded. Leveraging its modular expert architecture, RingMoE can be decomposed into modality-specific sub-models (e.g., RingMoE-Opt, RingMoE-MS, RingMoE-SARL1, RingMoE-SARL2). Each sub-model retains only its relevant modal-specialized, collaborative, and shared experts, rather than the full set across all modalities. This targeted expert retention reduces the parameter count from 14.7 billion to 6.5 billion, significantly lowering computational demands while maintaining performance (see Appendix Fig.A3.a).
To enable efficient deployment in resource-constrained environments, we introduce tailored expert pruning strategies to further compress the model while preserving its effectiveness (see Fig.7). By selectively removing underutilized experts, the model undergoes further pruning (RingMoE-EP) while retaining essential components, such as the multi-head attention layers. For instance, RingMoE-Opt achieves a parameter reduction to 4.3 billion. Additionally, we introduce expert knowledge aggregation, transforming the sparse model into dense configurations by integrating expert knowledge through summing, averaging, or compressing. This results in RingMoE-KS, RingMoE-KA, and RingMoE-KC, each with 1 billion parameters, maintaining accuracy while significantly improving deployment flexibility. Detailed methodologies for pruning are as follows:
@@ 3.4.1 Sparse Expert Pruning Schemes
To maximize knowledge retention in RingMoE, we initially retain all experts but also explore pruning unused or infrequently utilized ones. Specifically, for cooperative experts designed to capture inter-modal commonalities, we observe that despite implementing load-balancing loss to distribute the workload evenly across experts, certain experts are preferentially selected for tokens from specific modalities. This results in varying usage frequencies across experts when handling single-modal inputs. Therefore, we calculate each cooperative expert’s usage frequency based on input modality and prune those with a frequency below a threshold $\varphi$, tailored to each modality:
@@
Here, $f\left(E_{k}^{C}|x_{m}\right)$ represents the activation frequency of the $k^{th}$ cooperative expert for tokens from modality $m$ during pre-training. By pruning cooperative experts with low activation frequencies, we reduce the model’s parameters without significantly affecting its performance. Notably, modal-specialized and the shared experts in the RMoE layer remain untouched, preserving their full capacity to handle modal-specific tasks and shared knowledge across modalities. As a result, the expert pruning strategy produces the compressed RingMoE-Opt model with 4.3 billion parameters.

### v2

@@ 1 Introduction
Foundation models have recently emerged as transformative in AI, leveraging self-supervised learning [5] to extract generalized, task-agnostic representations from large-scale unlabeled data. Compared to smaller, task-specific models tailored for individual scenarios, these “one-for-all” general foundation models demonstrate superior performance and generalization across diverse interpretation tasks, particularly in data-scarce environments. Notable examples such as ChatGPT [6] in NLP, SiMIM [7] and MAE [8] in CV, and emerging models [9, 10, 11] in specialized fields like medical imaging and meteorology have redefined benchmarks in their respective domains.
@@
(1) Unified Multi-modal Perception. Building a complete Earth observation system usually requires multiple sensors (e.g., Opt, MS, and SAR) working in tandem to capture Earth’s surface properties. However, most existing RSFMs only serve single or narrow modalities (see Tab.I), e.g. RingMo [15] focus on Opt and DINO-MM [38] on Opt and SAR. Indeed, such modality-specific tailored models hinder the enormous potential of exploring RS multi-modal data. Numerous works have revealed that although heterogeneous sensors have different imaging modes (active imaging for SAR and passive imaging for Opt/MS) and non-uniform technological regimes (high-resolution imaging, wide-area observation, and dynamic monitoring, etc), the information from different modalities can complement each other [39], e.g., the fusion of Opt and SAR can improve the imaging quality during nighttime and bad weather conditions [40]. At this point, combining these modalities to construct a unified RS cross-modal representation can fully unleash the potential of multi-modal data and facilitate a better understanding of RS images.
Meanwhile, most existing SAR-based RSFMs utilize SAR amplitude data (namely SAR-L2 in this paper) as a visual proxy, treating it similarly to grayscale images for contrastive or masked image modeling [25, 20, 23]. While this simplification facilitates model design, it discards the phase component of the SAR signal, resulting in the loss of structural and scattering clues essential to understanding radar imaging physics [41]. In contrast, complex-valued SAR-L1 data retains both amplitude and phase, enabling richer physical interpretations of the scene. Recent studies [41, 42, 43] demonstrate that incorporating phase information helps capture physically grounded clues, such as surface roughness, dielectric properties, and object geometry, that are inaccessible via amplitude alone. Fully exploiting SAR-L1’s unique characteristics is therefore crucial for building RSFMs that move beyond visual approximations to a more faithful physical understanding of radar data.
@@ 2.1 Remote Sensing Foundation Models (RSFMs)
Self-supervised learning [5, 8, 7] in the CV field has significantly contributed to the flourishing of RSFMs [26, 25, 36, 17, 16, 19, 37, 23, 15, 22, 38], which can leverage the vast amount of unlabeled RS images to efficiently capture the generic feature representation. Mainstream RSFMs mainly employ masked image modeling (MIM) and contrastive learning (CL) for pre-training. (1) MIM-based RSFM, which learns image representations by reconstructing masked patches. RingMo [15] optimizes the masking strategy in MIM to focus on dense and small objects. SpectralGPT [18] designs a 3D tensor-shaped spatial-spectral mask strategy to comprehensively capture locally spatial-spectral features and spectral sequential information. Scale-MAE [16] reconstructs low-frequency and high-frequency information of the masked image at different scales to improve multi-scale perception. (2) CL-based RSFM, which optimizes the model by closing the distance between positive samples and widening the distance between negative samples. GASSL [50] takes time-series views as positive sample pairs to neglect temporal variation and focus on more spatial variation. SeCo [51] constructs seasonal-based multi-augmentation contrast loss to perceive short- and long-term temporal changes.
@@ 2.3 Mixture-of-Experts (MoE)
Experience with deep learning suggests that increasing model capacity and dataset scale tends to improve performance [44]. However, it must be acknowledged that scaling models to mega-scale also requires extremely high computational costs. Mixture-of-Experts (MoE) architecture [53, 45, 47, 48] is a promising solution to keep the computational cost acceptable while scaling up the model. With the sparsity of the MoE, the NLP community has expanded language models to a considerable scale [48, 45, 53, 54]. Switch Transformer [48] optimizes the MoE routing pattern and builds a model with up to a trillion parameters. MoA [54] proposes a new architecture that combines multiple attention and the MoE technique. DeepseekMoE [45] constructs fine-grained expert segmentation and shared expert isolation to increase expert specialization and model parameter efficiency (see Appendix A for details on typical MoE architecture). In addition, the researchers have also applied the MoE technique to the CV field [33, 53, 55, 47], e.g, V-MoE first applies the MoE to the CV field and proposes a 15B parameter model, which shows the potential of scale vision models. $\mathbf{M}^{3}$ViT [55] and Mod-Squad [47] also design efficient multi-task learning models with the MoE technology. In this paper, we apply the MoE technique to RS and construct a 14.7B parameter RSFM. Notably, for maximizing the parameter potential, we propose modal-specialized, collaborative, and shared experts to rationally divide up the learning of RS multi-modal knowledge.
Figure 5: Overview of the proposed RingMoE framework. Given the multi-modal inputs, the random masking operation is performed to generate visible and masked patches, i.e. tokens. The RingMoE encoder incorporates the RMoE layer (i.e., a novel hierarchical Mixture-of-Experts structure) to replace all standard FFN layers in each stage, effectively capturing both inter-modal commonalities and intra-modal specializations in RS. Finally, the latent representations from different modalities are decoded by modal-specific decoders to reconstruct the original targets.
@@ 3 Proposed RingMoE
This section outlines the pre-training, architecture, efficient deployment, and pre-training dataset. Sec.3.1 provides an overview of the RingMoE pre-training process, followed by Sec.3.2 and Sec.3.3, which detail the feature encoder and modal-specific decoder. Sec.3.4 discusses the expert pruning for effective deployment on downstream tasks. Finally, Sec.3.5 introduces the multi-modal dataset for pre-training.
@@ 3.2 Joint Encoder across Modalities and Experts for Multi-Modal Representation
RingMoE utilizes Swin Transformer v2 [32] as its image encoder, replacing all standard Feed-Forward Network (FFN) layers with the MoE layer to enhance model capacity and improve adaptability to multi-modal RS data (detailed in Appendix B.2). As illustrated in Fig.5, the image tokens from different modalities $x_{m}$, are first transformed into feature embeddings by modal-specific linear projections $E_{e}$. Then, after embedding the position encoding $E_{pos}$, all the feature embeddings $z_{m}$ from different modalities are fed into the RingMoE $f_{\theta}$ to learn the latent representations $z_{m}$, i.e., $z_{m}=f_{\theta}\left(E_{e}x_{m}+E_{pos}\right)$. Notably, the encoder $f_{\theta}$ consists of $L$ blocks, each of which is built from the Multi-head attention (MHA) and RMoE Layer.
@@ 3.2.1 RMoE Layer
To further enhance model capacity and fully exploit the potential of RSFMs, we propose replacing each FFN layer in the Swin Transformer with the MoE structure (see Appendix A for typical MoE implementations), where multiple parallel experts are dynamically activated through a routing network that assigns each token to its most relevant experts based on learned gating scores. However, singularly sharing a group of experts across different RS modalities may lead to interference between unrelated or even conflicting modalities, causing confusion or degradation in expert knowledge. In other words, while sharing experts across certain modalities can be beneficial, it can also be detrimental when applied to modalities with distinct imaging characteristics, such as Opt and complex-valued SAR-L1 data. This challenge highlights the inherent difficulty of training a unified multi-modal RSFM.
@@
Shared expert. The shared expert consolidates global knowledge across modalities, bypassing the complexity of a routing network by directly processing feature tokens from all inputs. This design reduces reliance on additional parameters in the modal-specialized and collaborative experts, enhancing efficiency while maintaining strong performance. By integrating cross-modal information, the shared expert reduces redundancy and ensures robust generalization.
@@ 3.2.2 Optimization of the RMoE layer
$f_{k}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left(\mathrm{Token}\ x_{m;i}\ \mathrm{selects}\ \mathrm{Expert}\ k\right)$ (9)
@@
Where $N_{C}$ denotes the number of cooperative experts. $f_{k}$ denotes the probability of tokens dispatched to expert $k$ (see Eq.9), where $N$ denotes the number of the input tokens, $\mathbb{I}(\cdot)$ is used to determine whether the $i^{th}$ input token $x_{m}$ is assigned to expert $k$ after passing through the routing network $G^{C}$ of the cooperating experts. $H\left(\cdot\right)$ denote the routing network $G^{C}\left(\cdot\right)$ without top-K function.
@@ 3.3 Modal-Specific Decoders for Physics-Informed Self-Supervised Target Reconstruction
For Opt, MS, and SAR-L2 images, we reconstruct original pixel values to preserve spatial and spectral integrity. For fully polarimetric SAR-L1 data, which consists of four polarization modes with complex-valued representations (each containing real and imaginary components), we adopt a physics-informed reconstruction strategy. Instead of reconstructing raw pixel values, we leverage the principle of power conservation before and after polarization decomposition [49], reconstructing the image power as: $t_{SAR_{L1}}=\mathrm{Power}\left(I_{SAR_{L1}}\right)=\left|I_{SAR_{L1}}^{HH}\right|^{2}+\left|I_{SAR_{L1}}^{HV}\right|^{2}+\left|I_{SAR_{L1}}^{VH}\right|^{2}+\left|I_{SAR_{L1}}^{VV}\right|^{2}$. Here, HH, HV, VH, and VV represent the four polarization modes. This approach ensures the reconstruction task remains physically meaningful, enhancing the model’s ability to capture SAR polarimetric characteristics.
@@
Where $\hat{x}_{m}$ and $t_{m}$ denote the reconstructed and original targets from modal $m$, respectively. $M\left(\cdot\right)$ denotes the set of masked pixels and $\Omega\left(\cdot\right)$ denotes the number of elements.
@@
Where $\alpha$ balances the contribution of two losses.
Figure 7: Details of Expert pruning. After pre-training, RingMoE supports modular decomposition into single-modal or multi-modal sub-models of varying scales, enabling flexible adaptation to downstream tasks. RingMoE can be compressed into modality-specific (e.g., 6.5B) or lightweight (e.g., 1B) models through expert pruning, retaining adaptability for resource-constrained environments with minimal performance degradation.
@@ 3.4 Dynamic Expert Pruning for Efficient Deployment across Various Computational Resources
Following large-scale pre-training, RingMoE ($f_{\theta}$) serves as a universal image encoder for downstream tasks, while the modal-specific decoders ($g_{\phi}$) are discarded. Leveraging its modular expert architecture, RingMoE can be decomposed into modality-specific sub-models (e.g., RingMoE-Opt, RingMoE-MS, RingMoE-SARL1, RingMoE-SARL2). Each sub-model retains only its relevant modal-specialized, collaborative, and shared experts, rather than the full set across all modalities. This targeted expert retention reduces the parameter count from 14.7 billion to 6.5 billion, significantly lowering computational demands while maintaining performance (see Appendix Fig.3.a).
To enable efficient deployment in resource-constrained environments, we introduce tailored expert pruning strategies to further compress the model while preserving its effectiveness (see Fig.7). By selectively removing underutilized experts, the model undergoes further pruning (RingMoE-EP) while retaining essential components, such as the multi-head attention layers. For instance, RingMoE-Opt achieves a parameter reduction to 4.3 billion. Additionally, we introduce expert knowledge aggregation, transforming the sparse model into dense configurations by integrating expert knowledge through summing, averaging, or compressing. This results in RingMoE-KS, RingMoE-KA, and RingMoE-KC, each with 1 billion parameters, maintaining accuracy while significantly improving deployment flexibility. Methodologies for pruning are as follows (see Appendix B.4 for more details):
@@ 3.4.1 Sparse Expert Pruning Schemes
To maximize knowledge retention in RingMoE, we initially retain all experts but also explore pruning unused or infrequently utilized ones. Specifically, for cooperative experts designed to capture inter-modal commonalities, we observe that despite implementing load-balancing loss to distribute the workload evenly across experts, certain experts are preferentially selected for tokens from specific modalities. This results in varying usage frequencies across experts when handling single-modal inputs.
To quantify this phenomenon, we compute the activation frequency $f(E_{k}^{C}|x_{m})$ of each collaborative expert $E_{k}^{C}$ on tokens from modality m, and define the retained expert set:
@@
Here, $\varphi$ denotes a modality-specific pruning threshold. To eliminate underutilized collaborative experts, we employ a frequency-aware expert pruning strategy (see Appendix B.4 for details). For each RMoE layer and modality $m$, we retain only the top 25% of the most frequently activated collaborative experts, where $\varphi$ corresponds to the 75th percentile of the activation frequency distribution. This adaptive thresholding effectively prunes redundant experts while preserving those critical for modality representation, leading to a substantial reduction in model size with minimal performance loss. Importantly, the modal-specialized and shared experts remain intact, ensuring full capacity for capturing both modality-specific and cross-modal knowledge. Consequently, the pruning strategy yields a 4.3B-parameter RingMoE-Opt model that achieves an optimal balance between efficiency and accuracy.